OpenStudy (anonymous):

The following is a graph of y = sin2 x + cos2 x + k . Find the value of k. the graph is a horizontal line y=6 a)1 b)6 c)4 d)5 answer is d please explain how to get to the answer, thank you

OpenStudy (dumbcow):

sin^2 +cos^2 = 1 , this is pythagorean trig identity y = 1+k = 6 k = 6 -1 = 5

OpenStudy (anonymous):

oh! I get it, thank you!